OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (greencat):

y=mx+b

OpenStudy (anonymous):

thats not a choice

OpenStudy (greencat):

no that is the beginning of solving it

OpenStudy (rational):

\[y-b = b(x-a)\]

OpenStudy (godlovesme):

U have choices? @izzyvelazco

OpenStudy (greencat):

yes i realized i was using wrong equation

OpenStudy (rational):

it is weird to use the letter "b" for slope lol

OpenStudy (greencat):

@rational don't give direct answers

OpenStudy (greencat):

just saying...

OpenStudy (anonymous):

y = −bx − b(1 − a) y = bx − b(a − 1) y= 1/b x +b (a-1) y= -1/b x+ b (1-a)

OpenStudy (anonymous):

nevermind this is stupid pellet flutter this site

OpenStudy (rational):

that was the point slope form of line @GreenCat , not the answer

OpenStudy (rational):

\[y-b = b(x-a)\] add \(b\) to both sides and simply eyeball the options @izzyvelazco

OpenStudy (greencat):

I knw but it was pretty close

OpenStudy (rational):

hmm